OpenStudy (anonymous):

I am trying to understand the whole idea of finding zero vectors, How do you find them?

OpenStudy (amistre64):

row reduction ....

OpenStudy (amistre64):

can you be more specific? im assuming this is a linear algebra concept, but i cant be sure

OpenStudy (anonymous):

yes, its linear algebra, but in vector spaces

OpenStudy (amistre64):

do you have a problem that can be used as an example?

OpenStudy (anonymous):

ok, let me type it now

OpenStudy (anonymous):

Let V = { * }, with vector addition +: V x V -->V defined by * +* = * and k* =* for k is an element of real numbers R. Check that the 10 axioms for vector spaces hold, and find the zero vector. Find -*

OpenStudy (anonymous):

please treat * as a big star

OpenStudy (amistre64):

so, it looks like the elements of the set consist of 1 thing: * addition (or the operation on the set) is such that *+*=*, so we have closure. And it seems like a defined property is given also as k* = * for some arbitrary real constant scalar.

OpenStudy (anonymous):

yes, it is closed under addition

OpenStudy (amistre64):

hmm, well the operation is also defined as its identity: e+* = * when e=* inverse; * + (-*) = e * + (-*) = * so it is also its own inverse .... i do not see a way to generally define Ax = 0 since k*=* and there is no 0 to be had.

OpenStudy (anonymous):

I don't know what to say man because I have no clue man

OpenStudy (amistre64):

does it say that this IS a vector space? or are you trying to prove that it is or is not a vector space?

OpenStudy (anonymous):

ok, i want u to help me may be with an example, on how to find zero vector of space, is it doable?

OpenStudy (amistre64):

depends on how much a recall :) but i can try

OpenStudy (anonymous):

please

OpenStudy (amistre64):

the zero vector can be determined, given a rule defining an operation, if the operation will produce 0

OpenStudy (amistre64):

if we define a+a = a, and ka = a then there is no way to really express a zero vector unless we assume a=0 to start with.

OpenStudy (anonymous):

and say a.0 =0+0 =0, that's where it hits me the most. like does the 0+0 come from

OpenStudy (amistre64):

it might help if you refresh my memory as to the 10 axioms of a vector space. I recall some things equating inverse with zero vectors, but its a bit fuzzy at the moment

OpenStudy (amistre64):

A real vector space is a set X with a special element 0, and three operations: Addition: Given two elements x, y in X, one can form the sum x+y, which is also an element of X. Inverse: Given an element x in X, one can form the inverse -x, which is also an element of X. Scalar multiplication: Given an element x in X and a real number c, one can form the product cx, which is also an element of X. These operations must satisfy the following axioms: Additive axioms. For every x,y,z in X, we have x+y = y+x. (commutative) (x+y)+z = x+(y+z). (associative) 0+x = x+0 = x. (identity) (-x) + x = x + (-x) = 0. (inverse) Multiplicative axioms. For every x in X and real numbers c,d, we have 0x = 0 (zero vector) 1x = x (identity) (cd)x = c(dx) (assiciative) Distributive axioms. For every x,y in X and real numbers c,d, we have c(x+y) = cx + cy. (c+d)x = cx +dx.

OpenStudy (amistre64):

http://www.math.uconn.edu/~troby/Math2210F13/LT/sec4_1.pdf here is another site that simply says: there is an element (the zero vector) such that * + z = *; in this case i would then say the zero vector, z=*

OpenStudy (anonymous):

yes, it is the zero vector that I seek to understand.How do you apply it inorder to find zero vector?

OpenStudy (amistre64):

for vector addition, the "zero vector" is the identity element. it just seems odd for a 1 element set.

OpenStudy (amistre64):

if we simply rename the information as: Let V = { 0 }, with vector addition +: V x V -->V defined by 0 +0 = 0 and k0 =0 for k in R. Check that the 10 axioms for vector spaces hold, and find the zero vector. Find -* it is isomorphic to the original information and easier to "see"

OpenStudy (anonymous):

yeah. this stuff requires wide reading to get the head and the tail, but i really appreciate the website. I am checking it now. what u r saying now adds more sense to the whole qquestion.

OpenStudy (amistre64):

1, closure; 0+0 = 0 is good 2, commutative; 0 + 0 = 0 + 0 is good 3, associative; 0 + (0 + 0) = 0 + 0 = 0 (0 + 0) + 0 = 0 + 0 = 0 is good 4, there is an e element in V such that 0 + e = 0 e = 0 is an element in V 5, inverse; there is an element i in V such that 0 + i = 0 i = 0 is an element in V 6, closure again k0 = 0 ... is good 7 and 8 are distributive laws: k(0+0) = k0 + k0 = 0 + 0, are fine 9 is associative again, c(k0) = c0 = 0 (ck)0 = 0 since ck is just some real value 10 indentity: 1(0) = 0 seems good

OpenStudy (amistre64):

if we change all the 0s in that to *s, all the properties still hold. It is just simpler to "see" with the element being a 0 :)

OpenStudy (anonymous):

thanx @amistre64 . really appreciate this

OpenStudy (amistre64):

youre welcome.